# Vision-driven Automated Mobile GUI Testing via Multimodal Large Language Model

### v1

@@ Abstract
With the advancement of software rendering techniques, GUI pages in mobile apps now encompass a wealth of visual information, where the visual semantics of each page contribute to the overall app logic, presenting new challenges to software testing. Despite the progress in automated Graphical User Interface (GUI) testing, the absence of testing oracles has constrained its efficacy to identify only crash bugs with evident abnormal signals. Nonetheless, there are still a considerable number of non-crash bugs, ranging from unexpected behaviors to misalignments, often evading detection by existing techniques. While these bugs can exhibit visual cues that serve as potential testing oracles, they often entail a sequence of screenshots, and detecting them necessitates an understanding of the operational logic among GUI page transitions, which is challenging traditional techniques. Considering the remarkable performance of Multimodal Large Language Models (MLLM) in visual and language understanding, this paper proposes a vision-driven automated GUI testing approach VisionDroid to detect non-crash functional bugs with MLLM. It begins by extracting GUI text information and aligning it with screenshots to form a vision prompt, enabling MLLM to understand GUI context. The function-aware explorer then employs MLLM for deeper and function-oriented GUI page exploration, while the logic-aware bug detector segments the entire exploration history into logically cohesive parts and prompts the MLLM for bug detection. We evaluate VisionDroid on three datasets and compare it with 10 baselines, demonstrating its excellent performance. The ablation study further proves the contribution of each module. Moreover, VisionDroid identifies 29 new bugs on Google Play, of which 19 have been confirmed and fixed.
@@
Automated GUI testing, Android app
@@ I Introduction
In an era where mobile applications are integral to everyday life, ensuring the quality and reliability of these applications is paramount [1]. The field of mobile app development is characterized by rapid iterations and the need for continuous integration and deployment, making the quality assurance process both critical and challenging. As one kind of acceptance testing, automated Graphical User Interface (GUI) testing [2, 3, 4, 5, 6, 7, 8, 9, 10] has become a cornerstone in this process, promising to streamline testing by systematically identifying issues that could degrade user experience.
Despite the advancements in automated testing techniques, the intricate nature of mobile interfaces, coupled with the variability of user interactions, often necessitates the human testers as the last line of defense in ensuring app quality. Besides severe app crashes with obvious oracle signals, many bugs present visual cues such as misalignments, unintended element overlaps, or deviations from expected results which current automated systems frequently overlook [11]. As demonstrated in Fig. 1, the first sub-figure exposes a bug where the total income (5,030) does not match the summarization of the detailed records. Additionally, the four sub-figures together unveil a bug related to the failure of adding a record, as the income-expense overview in the fourth sub-figure should display five records after adding a new one. Our empirical study in Section II identifies nearly nine categories of such bugs, indicating the diversity of these non-crash bugs with visual cues. Although these bugs exhibit visual cues that make visual detection possible, they often entail a sequence of screenshots (as demonstrated in Fig. 1), requiring the understanding of the overall operational logic among these GUI page transitions. Only human testers with both sharp eyes and knowledge from extensive training can identify those bugs.
@@
However, human testing is time-consuming and suffers from inherent scalability bugs. Although researchers begin to explore detecting non-crash or functional bugs with visual information, most of them are typically designed to identify specific types of bugs, such as display issues [12, 13], scaling problems [14, 15], game GUI glitches [16] or WebView bugs [17]. What makes it worse is that all of them require either a set of rules that are easy to be out of date, or a large amount of training dataset which takes much human effort for annotation. This specialization leaves a gap in the testing landscape for a more generalized, lightweight and intelligent approach that can adapt to the diverse challenges of mobile app testing especially for non-crash bugs.
Thanks to the emerging Multimodal Large Language Model (MLLM) [18], it is trained on the ultra-large-scale corpus and potentially able to understand the GUI pages as well as the information changes along with the page transition. Inspired by this, we propose a vision-driven automated GUI testing approach with MLLM. Nevertheless, three challenges need to be addressed. First, solely feeding the raw screenshot would be far less effective in letting the MLLM understand the GUI context, therefore how to align other sources of information and produce an effective vision prompt. Second, the triggering of bugs largely depends on whether and how the corresponding GUI page can be reached, therefore how to guide the MLLM to plan the exploration to delve deeper, covering diverse scenarios, and focusing on the app’s functionalities. Third, detecting the functional bugs requires tremendous domain knowledge and MLLM might lack it, therefore how to provide accurate information to facilitate MLLM making decisions.
To tackle these challenges, we propose VisionDroid, a vison-driven approach for detecting non-crash functional bugs with MLLM. It begins by extracting GUI text information and aligning it with GUI screenshots to form a vision prompt, enabling MLLM to understand GUI context. Second, the function-aware explorer utilizes the MLLM for GUI page exploration. It maintains a testing history comprising both the higher-level catalogue of testing progress and a detailed snapshot of each performed action, to enhance the exploration to be deeper and function-oriented. Third, the logic-aware bug detector breaks the entire exploration sequence into logically cohesive parts that mirror the app’s logical operational sequences. Based on them, the MLLM detects the bugs potentially by checking whether the GUI information changes between page transitions align with the processing logic. To boost the performance, we leverage in-context learning schema to provide useful examples when querying the MLLM, from online bug reports and history running records.
@@
To evaluate the effectiveness of VisionDroid, we experiment on three datasets. Compared with 10 common-used and state-of-the-art baselines, VisionDroid can achieve more than 8%-230% and 27%-43% boost in average recall and precision compared with the best baseline, resulting in 50%-76% precision and 42%-64% recall. To further understand the role of each sub-module of the approach, we conduct ablation experiments to further demonstrate its effectiveness. We also evaluate the usefulness of VisionDroid by detecting unseen non-crash bugs in real-world apps from Google Play. Among 105 apps, VisionDroid detects 29 new crash bugs with 19 of them being confirmed and fixed by developers, while the remaining are still pending.
@@
  • Vision. The pioneering work to formulate the non-crash functional bugs detection problem to an interactive multimodal question & answering task, enabling the MLLM to identify bugs by comprehending GUI visual information and functional logic.
@@
  • Technique. A vision-driven automated GUI testing approach VisionDroid [^1] We release the source code, dataset, and detailed experimental results on our website https://github.com/testtestA6/VisionDroid. which designs a function-aware explorer and logic-aware bug detector for detecting non-crash functional bugs via MLLM.
@@
  • Evaluation. Effectiveness and usefulness evaluation of the VisionDroid in the real-world apps with practical bugs detected (Section V-A and V-C).
II motivational study
To better understand the non-crash functional bugs in real-world practice, we carry out a pilot study to examine their prevalence. It also explores what categories of non-crash functional bugs, to facilitate the design of our approach.
II-A Data Collection
Our experimental dataset is collected from one of the largest open-source Android app hosting websites (GitHub) [19], which has a great number of Android bug reports. We employ the following criteria for app selection: more than 1K downloads in Google Play (popular), a public issue tracking system (traceable), and more than three years of development history (trustworthy). A total of 3,128 open-source Android apps with 57,326 bug reports are selected in this study. These apps belong to diversified categories such as business, communication, finance, etc. For the selected apps, we randomly select 6,000 bug reports involving 576 apps with detailed bug descriptions and GIFs/screenshots, and conduct manual analysis. Among them, 4,856 (81%) bug reports are related to the functional aspects of the app, while the rest are about the performance, security, etc. These functional bugs include 1,796 crash bugs (37%), and 3,060 non-crash functional bugs (63%). Since the crash bugs have been widely studied in previous work, we further investigate the non-crash functional bugs.
Figure 2: The categories of non-crash functional bugs
Figure 3: Overview of VisionDroid
II-B The Categories of Non-crash Functional Bugs
For the above-mentioned 3,060 non-crash functional bug reports, we manually classify them based on bug descriptions and reproduction steps, and derive the categories of non-crash functional bugs to attain a comprehensive understanding of them. Following the open coding protocol [20], two authors with more than three years of development and testing experience individually examine the issue description and reproduction steps. Then each annotator iteratively merges similar codes, and any disagreement of the categorization will be handed over to the third researcher for double-checking. Finally, we come out with a categorization of the non-crash functional bugs within (intra-page) and among the pages (inter-page), with details summarized in Fig. 2. We additionally provide examples for each category in our website1.
Intra-page non-crash functional bug. Intra-page non-crash functional bug denotes these bugs only related to a single GUI page, mainly involving the issues about information display as shown in Fig. 2.
Inter-page non-crash bug. Inter-page non-crash functional bug denotes these bugs related to multiple GUI pages, i.e., involving a testing sequence to trigger it, with details in Fig. 2. Among them, data operation failures mainly manifested as failures in adding, deleting, and modifying data, such as failed user information updates, errors in adding or removing shopping cart items, etc. Media control no-response mainly involves issues with video and audio playback, such as no response of playback control (failure of play, pause, stop functions). Numerical calculation errors mainly include mathematical errors, inaccurate statistical data, and cost calculation deviations. Setting and configuration failures mainly manifested as users being unable to save personalized settings or configurations, including ineffective or reset settings within the app (such as notification switches and theme changes). Navigation and link errors mainly manifested as page redirection errors and confusion in navigation logic between functions.
Findings. The above analysis confirms the popularity of non-crash bugs in mobile applications, implying the need for comprehensive testing. Meanwhile, considering the diversity of application functionality and bug manifestation, manually designing bug detection rules and building a complete set of defect methods require significant effort, and the automated technique is highly demanded. Despite their complexity, we have observed that these non-crash functional bugs typically manifest visual cues, either within a single GUI page or across a sequence of GUI pages. This motivates us to apply visual-driven techniques for bug detection.
III Approach
This paper proposes VisionDroid, a vision-driven automated GUI testing approach for detecting non-crash functional bugs with the Multimodal Large Language Model (MLLM). As shown in Fig. 3, it begins by extracting GUI text information, and aligning it with GUI screenshots, forming the vision prompt for GUI context perception. Second, the function-aware explorer utilizes the MLLM for guiding the exploration to delve deeper, encompassing more diverse pages, and focusing on the app’s functionalities. This is facilitated through maintaining the testing history comprising both the higher-level catalog of testing progress and the detailed snapshot of each performed action. Third, the logic-aware bug detector breaks the entire exploration sequence into locally cohesive parts that mirror the app’s logical operational sequences. Based on them, the MLLM detects the bugs potentially by checking whether the GUI information changes between page transitions align with the processing logic. To further boost the performance, we leverage the in-context learning schema to provide useful examples when querying the MLLM, from online issue reports and history running records.
III-A Alignment between Text and Image
@@
Existing automated GUI testing approaches typically rely on view hierarchy files to obtain the attribute information (text, coordinates, actions, etc.) of app widgets to capture the current context and determine which widgets should be acted upon. This could be less efficient since the conversion from screenshot image to text could potentially lose information about the structure semantics of the GUI. Thanks to the MLLM, we can directly feed the screenshot image as input. Yet solely based on the screenshot also has its limitations since the image tends to utilize icons and extremely concise text for conveying information, which can also hinder the MLLM’s perceptions. Therefore, we design the visual and text alignment method, aiming at enhancing the MLLM’s understanding of the semantics of the GUI page.
III-A1 GUI Text Information Extraction
@@
We extract three types of information, i.e., the app, the current GUI page, and all its contained widgets. The app information is extracted from the AndroidManifest.xml file, while the other two types of information are extracted from the view hierarchy file of the current GUI page, which can be obtained by UIAutomator [21]. We first extract the app information including the app name, and the activity names in this app , which can facilitate the MLLM to gain a general perspective of the functions. Then, we extract the page information from the view hierarchy file and extract the widgets information including “text” field, “hint-text” field, “resource-id” field, “class” field, “clickable” field, “long-clickable” field, “checkable” field, “scroll” field and “bounds” field, which facilitates the MLLM in suggesting actions related to these widgets.
@@
III-A2 Text-Image Alignment for Visual Prompt
We annotate the GUI pages to guide the LLM’s attention towards the actionable parts (i.e., widgets). as shown in the right part of Fig. 4. Following existing studies [6, 7, 8, 9, 10], we support five types of basic operational actions, i.e., click, input, long-click, check, and scroll. We draw bounding boxes for each of these operational widgets, which are based on the extracted widget coordinates from the property of the “bounds” field, and use different colors indicating different types. In detail, if the “clickable” property of a widget is “true”, we draw a red bounding box to indicate it can be clicked, and if the “class” property is “EditText”, we draw a blue bounding box to indicate it needs input text. Similarly, we respectively use the purple, pink and yellow bounding boxes to indicate the widget with “long-clickable”, “checkable”, and “scrollable” properties as “true”. Inspired by the screen reader [22], we reconstruct the coordinate system in order from top to bottom and from left to right, and label each action widget on the image using Arabic numerals starting from 1. To mitigate the crowded appearance resulting from the abundance of components on a GUI page, we only mark the first widget of each row and let the MLLM infer others as shown in the prompt on the right side of Fig 4.
@@ III-A2 Text-Image Alignment for Visual Prompt
We then associate each annotated widget in the screenshot with the extracted GUI text information. Note that, in the “prompts related to GUI text and visual information” (details in Section III-B2 and Fig. 4), we would not provide the detailed alignment information considering the input length constraints of the MLLM. Instead, this alignment information will be used in “prompts related to action & feedback” (details in Section III-B2 and Fig. 5) to help verify whether the MLLM is suggesting interactive widgets in accordance with the correct logical inference process, akin to an explicit chain-of-thought mechanism. In detail, when querying the general action or text input, we not only request the MLLM to provide the widget ID from the screenshot, but we also ask for the corresponding widget text. When these two sources of information do not align, we infer potential errors in the MLLM’s decision-making process and prompt it to re-try.
III-B Function-aware Explorer
The non-crash functional bugs are often linked to the functionalities of the app, which often require lengthy operational sequences to trigger. Consequently, we design the function-aware explorer to guide the exploration of the app’s functionalities, while also encouraging deeper dives and broader coverage of GUI pages.

### v2

# Seeing is Believing: Vision-driven Non-crash Functional Bug Detection for Mobile Apps

@@ Abstract
Mobile app GUI (Graphical User Interface) pages now contain rich visual information, with the visual semantics of each page helping users understand the application logic. However, these complex visual and functional logic present new challenges to software testing. Existing automated GUI testing methods, constrained by the lack of reliable testing oracles, are limited to detecting crash bugs with obvious abnormal signals. Consequently, many non-crash functional bugs, ranging from unexpected behaviors to logical errors, often evade detection by current techniques. While these non-crash functional bugs can exhibit visual cues that serve as potential testing oracles, they often entail a sequence of screenshots, and detecting them necessitates an understanding of the operational logic among GUI page transitions, which is challenging traditional techniques. Considering the remarkable performance of Multimodal Large Language Models (MLLM) in visual and language understanding, this paper proposes Trident, a novel vision-driven, multi-agent collaborative automated GUI testing approach for detecting non-crash functional bugs. It comprises three agents: Explorer, Monitor, and Detector, to guide the exploration, oversee the testing progress, and spot issues. We also address several challenges, i.e., align visual and textual information for MLLM input, achieve functionality-oriented exploration, and infer test oracles for non-crash bugs, to enhance the performance of functionality bug detection. We evaluate Trident on 590 non-crash bugs and compare it with 12 baselines, it can achieve more than 14%-112% and 108%-147% boost in average recall and precision compared with the best baseline. The ablation study further proves the contribution of each module. Moreover, Trident identifies 43 new bugs on Google Play, of which 31 have been fixed.
@@
Multimodal Large Language Model, Android app, Non-crash bugs
@@ I Introduction
With the proliferation of mobile applications, their impact on everyday life and commercial operations has grown significantly. To meet the increasing demands for advanced functionalities, app features have become more complex, inevitably leading to the introduction of bugs. Existing research primarily focuses on crash bugs due to their clear test oracles  [1, 2, 3, 4, 5, 6, 7, 8, 9]. However, there is a substantial number of non-crash functional bugs that remain underexplored. According to our statistics on over 100,000 GitHub issue reports from 2021 to 2024, approximately 61.3% describe non-crash functional bugs, with 47.7% of them tagged as “important”. These non-crash functional bugs can significantly impair functionality and even lead to disastrous consequences in various sectors and everyday routines. For instance, on February 29, 2024, a software payment functionality bug caused widespread malfunctions in self-service payment systems at gas stations across New Zealand, resulting in severe repercussions [10].
On the other hand, we have observed that many non-crash functional bugs reveal clear visual cues. While some of these cues are obvious, such as component occlusions and text overlaps as investigated in previous studies [11, 12], more often they involve issues with the visual logic of the page. For instance, in the first bug shown in Fig. 1, the page is supposed to display items sorted by the lowest price, but categories with lower prices appear at the end instead. Similarly, in the third bug, clicking “wish” fails to display the corresponding item on the wishlist page. These bugs necessitate not only a semantic understanding of the GUI page content but also a comprehension of the overall operational logic underlying these GUI page transitions, making automated detection significantly challenging. Only human testers with both sharp eyes and knowledge from extensive training can identify those bugs.
@@
However, human testing is time-consuming and suffers from inherent scalability and high-cost issues. Although researchers begin to explore detecting non-crash or functional bugs with visual information, most of them are typically designed to identify specific types of bugs (e.g., display [11], game glitches [13]). What makes it worse is that all of them require either a set of rules that are easy to be out of date, or a large amount of training dataset which takes much human effort for annotation. This specialization leaves a gap in the testing landscape for a more generalized, lightweight and intelligent approach that can adapt to the diverse challenges of mobile app testing especially for non-crash bugs.
Thanks to the emerging Multimodal Large Language Model (MLLM) [14], it is trained on the ultra-large-scale corpus and potentially able to understand the GUI pages as well as the information changes along with the page transition. Inspired by this, we propose Trident, a novel vision-driven, multi-agent collaborative automated GUI testing approach for detecting non-crash functional bugs. It comprises three agents: Explorer, Monitor, and Detector. The Explorer Agent navigates through the app, captures view hierarchies and screenshots, and guides the exploration towards diverse GUI pages while focusing on the app’s functionalities. The Monitor Agent supervises the testing process, records the exploration history, and triggers the detector agent at the appropriate time. The Detector Agent identifies potential functional bugs by examining whether there are any issues in the logical transitions that occur during GUI page changes.
While the design of Trident’s three agents is intuitive, improving their effectiveness in revealing functional bugs presents several challenges. The first challenge is aligning visual and textual information for MLLM input. We develop a screenshot annotation method to guide the MLLM to better understand the GUI page, along with an alignment method that integrates text properties with visual context to enable the MLLM to capture comprehensive information. The second challenge involves achieving functionality-oriented exploration. We design a method to infer and abstract the tested functionality from detailed exploration sequences, and use the abstracted history when communicating with MLLM to avoid exceeding token limits and enhance the detection of functional bugs. The third challenge is inferring test oracles for non-crash bugs. We design a mechanism to determine when to trigger the detector agent, and develop the functionality-aware chain-of-thought method to enable the MLLM to first explicitly infer oracles and then detect functional bugs based on these inferences, potentially reducing hallucinations in MLLM’s decision-making process.
@@
We evaluate the effectiveness of Trident on 590 non-crash bugs. Compared with 12 common-used and state-of-the-art baselines, Trident can achieve more than 14%-112% and 108%-147% boost in average recall and precision compared with the best baseline, resulting in 50%-72% precision and 42%-52% recall. To further understand the role of each sub-module of the approach, we conduct ablation experiments to further demonstrate its effectiveness. We also evaluate the usefulness of Trident by detecting unseen non-crash bugs in real-world apps from Google Play. Among 187 apps, Trident detects 43 new crash bugs with 31 of them being fixed and 12 of them being confirmed by developers.
@@
  • The pioneering work to formulate the non-crash functional bugs detection problem to an interactive multimodal question & answering task, enabling the MLLM to identify bugs by comprehending GUI visual information and functional logic.
@@
  • A vision-driven automated GUI testing approach Trident [^1] We release the source code, dataset, and experimental results on our website https://github.com/testtest2024-art/Trident. which includes a vision-driven explorer, process-aware monitor, and function-aware detector to identify non-crash functional bugs through MLLM.
@@
  • Effectiveness and usefulness evaluation of the Trident in real-world apps with practical bugs detected and confirmed (Section IV-A and IV-C).
Figure 2: Overview of Trident
II Approach
This paper proposes Trident, a novel vision-driven, multi-agent collaborative automated GUI testing approach for detecting non-crash functional bugs. Inspired by manual GUI testing, where human testers navigate through the app, observe the testing process, and identify bugs through visual and logical analysis, Trident comprises three agents: Explorer, Monitor, and Detector, as demonstrated in Fig. 2.
The Explorer Agent navigates through the app, captures view hierarchies and screenshots, and guides the exploration towards diverse GUI pages while focusing on the app’s functionalities. The Monitor Agent supervises the testing process, records the exploration history, and triggers the detector agent at the appropriate time. The Detector Agent identifies potential functional bugs by examining whether there are any issues in the logical transitions that occur during GUI page changes.
While the design of Trident’s three agents is intuitive, how to make them more effective in revealing functional bugs is difficult. Below, we outline three specific challenges.
Challenge 1: Aligning visual and text for MLLM input. Although MLLMs can directly process images and plan explorations based on screenshots, their outputs are limited to text. This means that MLLMs cannot directly mark the next actionable widgets on a screenshot; they must rely on textual descriptions. To cope with this, we propose an alignment method that integrates text properties with visual context (see Section II-A). This allows the MLLM to have a more comprehensive understanding of the GUI, thereby better guiding subsequent exploration. Furthermore, screenshots often contain numerous elements. If fed directly to the MLLM, it may overlook critical details. To mitigate this, we design a screenshot annotation method, guiding the MLLM to pay attention to different types of actionable widgets, and also resolve the issue of overlapping annotations caused by the density of widgets on the GUI page.
Challenge 2: Functionality-oriented exploration. During app testing, a sequence of actions and accompanying page transitions is generated. These actions focus on the operational level of the app, making it difficult for LLM to plan exploration paths from a functional perspective, thus impacting the detection of functional bugs. To address this, we develop a method to infer and abstract the current functionality from detailed exploration sequences (in Section II-B). It helps in two ways: 1) avoids exceeding token limits when interacting with the LLM; 2) enables the exploration more focusing on the functionality aspect for identifying functional bugs.
Challenge 3: Inferring test oracle. Accurately obtaining test oracles is a time-consuming and labor-intensive task, which has long been a limiting factor in the development of automated testing [15, 16, 6, 17, 7]. In GUI testing, it requires determining what the current GUI page should be like after each performed action. This not only involves information from a single page (e.g., missing text) but also considers a series of previous pages and actions (e.g., an item added not appearing on the list page), which is quite challenging. To address this, we let the Monitor Agent trigger the Detector Agent at the end of each functionality exploration (see Section II-B). This allows the Detector Agent to assess functionality by analyzing a manageable sequence of actions and states, thus focusing on the functionality and providing accurate inferences. Additionally, we develop the functionality-aware Chain-of-Thought (COT) to enable the MLLM to first explicitly infer oracles and then detect functional bugs based on these inferences, potentially reducing the hallucinations in MLLM’s decision-making process (in Section II-C1).
II-A Vision-driven Explorer Agent
@@
The Explorer Agent serves as the foundation of Trident, which visually inspects a GUI page, understands its semantics, and conducts the exploration. Before the MLLM, existing automated GUI testing approaches [16, 15, 18, 6, 17] typically relied on view hierarchy files to extract attribute information (text, coordinates, actions, etc.) of app widgets to capture the current context and determine which widgets should be acted upon. This can be inefficient as it may lose the structural semantics of the GUI when converting screenshots to text. Thanks to the MLLM, it allows us to use screenshot images as input directly. However, relying solely on images can also be limited due to the concise nature of GUI text and icons. Therefore, we design a method to align visual and textual information to enhance the MLLM’s understanding of the GUI page semantics.
II-A1 Alignment between Text and Image
@@
For the text, we extract three types of information: app details, current GUI page context, and the contained widgets. The app information is obtained from the AndroidManifest.xml file, while page and widget information are extracted from the view hierarchy file of the current GUI page using UIAutomator [19]. Initially, we gather information such as the {App name}, {Activities} and {Activity name} to provide the MLLM with a general understanding of the app’s functionalities. Subsequently, we extract page information and widget information including fields such as “text”, “hint-text”, “resource-id”, “class”, “clickable”, “long-clickable”, “checkable”, “scroll”, and “bounds”, which facilitate the MLLM in suggesting appropriate actions related to these widgets.
@@
For the image, we design a screenshot annotation method to guide the MLLM’s attention towards actionable widgets, as illustrated in Fig. 3. Trident supports five basic operational actions: click, input, long-click, check, and scroll, while the less standardized actions such as drag-and-drop or multi-finger gestures are excluded, following existing studies [5, 6, 7, 8, 9]. Bounding boxes are drawn around each actionable widget based on their coordinates from the “bounds” field, with different colors indicating different action types. For instance, a red bounding box indicates a “clickable” widget, a blue box indicates an “EditText” widget requiring text input, and other colors are used for long-clickable, checkable, and scrollable widgets. If a widget supports multiple actions, a green bounding box is used, with specific actions described in the accompanying text. Inspired by the screen reader [20], we label each widget with numbers from top to bottom and left to right, marking only the first widget of each row to reduce the overlapping annotations and letting the MLLM infer others. To help the MLLM better understand the annotation information on the image, we also provide a {legend of image} in the prompt as shown in Fig. 4.
Figure 4: Example of explorer prompt
@@
We associate each annotated widget in the screenshot with the extracted GUI text information. In prompts related to text and image information (details in Fig. 4), detailed alignment information is omitted due to input length constraints of the MLLM. Instead, this alignment is used to verify whether the MLLM suggests interactive widgets according to the correct logical process, akin to an explicit chain-of-thought mechanism. When querying general actions or text input, we request the widget ID from the screenshot and the corresponding widget text. If these sources do not align, we infer potential errors in the MLLM’s decision-making and prompt it to retry. To prevent continuous errors and infinite loops, we limit retries to three attempts before triggering a fallback strategy.
II-A2 Explorer Prompt
Fig. 4 demonstrates the explorer prompt. It is composed of the GUI information and screenshots as described in this section, and the testing history outlining the dynamic testing process provided by the Monitor agent (in Section II-B), as well as the explore optimization feedback from the Detector agent (in Section II-C). It queries the MLLM about the interactive widget and the action that needs to be taken.
II-B Process-aware Monitor Agent
The Monitor Agent acts as the central control unit of Trident, overseeing the test progress, abstracting the tested functionalities history for the Explorer Agent, and triggering the Detector Agent at appropriate times. It continuously captures the fine-grained exploration sequence of the app, including view hierarchies and screenshots of each explored GUI page, and the related actions on those pages. Based on this information, it generates two views of testing histories: a higher-level text view in the form of explored functionalities and a detailed snapshot of testing records in an image view. During the testing process, these views, which represent transformed testing histories, are provided to the Explorer Agent, which addresses the token limit for LLM’s inputs while also enhancing the understanding of the testing process. During the monitoring process, if the Monitor Agent determines that the current functionality is still under exploration, it instructs the Explorer Agent to continue. Once it identifies that a new functionality starts to test, the Monitor Agent pauses the Explorer Agent and triggers the Detector Agent to analyze potential bugs in the previously tested functionality.